OpenStudy (anonymous):

Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

b2 − 4ac 2x2 + 7x + 3 = 0 (7)^2-4(2)(3) 49-24 25 as 25 is a perfect square so the quadratic equation can be solved by factorisation

OpenStudy (anonymous):

b2-4ac, value here 25

OpenStudy (anonymous):

for ax^2+bx+c=0 b^2-4ac is the part that can tell 7^2-4*2*3 = 25